# A dynamic circuit for the honeycomb Floquet code

### v2

@@ VI Conclusion
We have demonstrated that our dynamic circuit increases the threshold and lowers the resource overhead of the Floquet code. Our work leaves open a few avenues for future exploration.
First, we have neglected to introduce boundaries in our circuits [15, 18], although we do not expect boundaries to significantly affect our results. We also have not simulated a stability experiment [28] to verify that the timelike error improves similar to the spacelike error rates. We also only simulated a simple depolarizing error model; it is likely that the relative advantage changes for other qubit-specific error models, and becomes even more significant for models that include leakage.
In addition, our simulations were decoded using minimum-weight perfect-matching (MWPM) [6]. It has been previously shown [14, 15] that a correlated decoder [29] can reduce the teraquop footprint of the Floquet code significantly compared to MWPM. It is an open question how an improved decoder would change the relative advantage of the dynamic circuit.
Finally, numerous generalizations of the Floquet code have been introduced that also reduce stabilizer measurements to a series of two-qubit gauge measurements [30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40]; it is an open question if these codes could also benefit from our dynamic approach. In particular, the square-octagon Floquet code [30] will not have a reduced spatial distance in the dynamic circuit, as the gauge operators do not lie along the logical operators; a circuit-level simulation of this code will likely lead to further improvements.
@@ Acknowledgements.
The ancilla-based simulations were performed using code included with [14]. The dynamic simulations also made substantial use of tools included in that code base.  Data availability–Stim circuits and simulation data are available at [41].
@@ Appendix A Noise Model
To be consistent with previous work [14, 15], we use the standard depolarizing noise model. The noise model consists of
@@
  • Initialization in $|0\rangle$ instead prepares $|1\rangle$ with probability $p$, and similar for |+ket|$\rangle$| + ⟩.
@@
Note that while our simulations of the dynamic and the standard circuits included idle errors, idle errors are significantly more damaging to the dynamic circuit, because once the standard circuit is compacted it has very few locations where idle errors can occur. If the dominant errors were only two-qubit gate errors, measurement errors, and reset errors, the advantage of the dynamic circuit would increase.
Appendix B Additional numerical results
Here, we include a plot of our extrapolation of logical error rate as a function of distance, used to compute the teraquop footprint for $p=10^{-3}$. We also include graphs of the separate horizontal and vertical logical error rates, rather than summed as in the main text. Note that the V logical error probability is the probability that the vertical logical operator is incorrect, i.e. the probability that an undetected horizontal logical operator has been applied, and similarly for the H logical error probability. The larger V logical error probability for the standard circuit was also seen in [15].
Figure 5: Additional numerical results, including our extrapolation to the teraquop regime for $p=10^{-3}$ and the logical error rates separated into horizontal and vertical error rates.

### v3

@@ VI Conclusion
We have demonstrated that our dynamic circuit increases the threshold and lowers the resource overhead of the honeycomb Floquet code. Our work leaves open a few avenues for future exploration.
First, we have neglected to introduce boundaries in our circuits [gidney2022benchmarking, haah2022boundaries], although we do not expect boundaries to significantly affect our results. We also have not simulated a stability experiment [gidney2022stability] to estimate the timelike error rates. We also only simulated a simple depolarizing error model; the advantage of the dynamic circuit will depend on the error model, and become larger for models that include leakage.
Finally, numerous generalizations of the honeycomb Floquet code have been introduced that also reduce stabilizer measurements to a series of two-qubit gauge measurements [paetznick2023performance, townsend2023floquetifying, davydova2023floquet, rodatz2024floquetifying, zhang2022x, dua2024engineering, de2024xyz, higgott2024constructions, fahimniya2023fault, alam2024dynamical, setiawan2024tailoring]; it is an open question if these codes could also benefit from our dynamic approach. In particular, the square-octagon Floquet code [paetznick2023performance] will not have a reduced spatial distance in the dynamic circuit, as the gauge operators do not lie along the logical operators; a circuit-level simulation of this code will likely show dramatic improvements.
@@ Acknowledgements.
The ancilla-based simulations were performed using code included with [gidney2021fault]. The dynamic simulations also made substantial use of tools included in that code base. Data availability–Stim circuits and simulation data are available at [zenodo].
@@ Appendix A Noise Model
To be consistent with previous work [gidney2021fault, gidney2022benchmarking], we use the standard depolarizing noise model:
@@
  • Initialization in $|0\rangle$ gets flipped with probability $p$, and similar for $|+\rangle$.
@@
Note that idle errors are significantly more damaging to the dynamic circuit, because once the standard circuit is compacted it has few idle locations. If idle errors were negligible, the advantage of the dynamic circuit would increase.
Appendix B Timelike distance
In the standard circuit, the $\otimes^{6}Z$ detector is formed as the product of gauge operators at times $XX(t)$, $YY(t)$, $XX(t+1)$, and $YY(t+1)$. Notably, this means that a detector spans four gauge measurements, and has a timelike extent of $4/3$ of a full round. In contrast, in the dynamic circuit, this detector is formed by multiplying $ZZ(t)$ and $YY(t+1)$ and has timelike extent of one full round (neither of these statements are obvious, and can best be understood by examining the circuit in Crumble).
We have numerically verified this behavior using STIM [gidney2021stim]. For fixed $d$, we generate circuits with a varying number $N_{m}$ of measurement rounds and produce their detector error models. We can then numerically find the minimum distance between detectors occuring around the beginning of the circuit (some fixed distance $\delta N_{m}$ from 0) and detectors occuring around the end of the circuit (some fixed distance $\delta N_{m}$ from $N_{m}$). The precise value of $N_{m}$ is not particularly important, as long as we keep it fixed as we scale $N_{m}$. We can then consider two quantities: (1) $d_{\text{graph}}$, the minimum number of graphlike errors it takes to connect the beginning detectors to end detectors (2) $d_{\text{hyper}}$, the minimum number of graphlike errors connecting the detectors if we decompose all hyperedges $h$ into complete graphs on the degree-$|h|$ vertices of $h$. These distances provide upper/lower bounds on the timelike distance $d_{t}$: $d_{\text{graph}}\geq d_{t}\geq d_{\text{hyper}}$. For the standard circuit we find $d_{\text{graph}}=d_{\text{hyper}}=3N_{m}/4$, so that we require $N_{m}=4d_{t}/3$ to reach distance $d_{t}$. For the dynamic circuit, we find $d_{\text{graph}}=2N_{m},d_{\text{hyper}}=N_{m}$, so we require at most $N_{m}=d_{t}$ to reach distance $d_{t}$ (but $N_{m}=d_{t}/2$ may be sufficient). Establishing the precise relationship between $d_{t}$ and $N_{m}$ for the dynamic circuit will require a stability simulation [gidney2022stability].
Appendix C Additional numerical results
Here, we include a plot of our extrapolation of logical error rate as a function of distance, used to compute the teraquop footprint for $p=10^{-3}$ as well as graphs of the separate horizontal and vertical logical error rates, rather than summed as in the main text. We present results from both MWPM decoding, as well as the more accurate but slower correlated matching decoding [fowler2013optimal]. The V logical error probability is the probability that the vertical logical operator is incorrect, i.e. the probability that an undetected horizontal logical operator has been applied, and similarly for the H logical error probability. The larger V logical error probability for the standard circuit was also seen in [gidney2022benchmarking]. The steeper slope of the V logical error probability in the dynamic circuit results from a complicated issue of error multiplicity that is difficult to fully account for. We note that all of these probabilities have a slope steeper than $p^{\lfloor d/2\rfloor}$ despite the circuits having a numerically-verified graphlike distance of $d$, due to higher-distance errors dominating at these values of $p$.
Figure 5: Our extrapolation to the teraquop regime for $p=10^{-3}$, and the logical error rates separated into horizontal and vertical error rates, for (left) MWPM and (right) correlated matching. Circuits simulated with STIM [gidney2021stim] and decoded with PyMatching [higgott2025sparse].